Required information The general ledger of Red Storm Cleaners at January 1, 2021, includes the following account balances: Accounts Debits Credits Cash $ 12,000 Accounts Receivable 6,400 Supplies 2,400 Equipment 18,000 Accumulated Depreciation $ 6,200 Salaries Payable 8,700 Common Stock 17,000 Retained Earnings 6,900 Totals $ 38,800 $ 38,800 The following is a summary of the transactions for the year: 1. March 12 Provide services to customers, $44,000, of which $19,400 is on account. 2. May 2 Collect on accounts receivable, $16,400. 3. June 30 Issue shares of common stock in exchange for $6,000 cash. 4. August 1 Pay salaries of $8,700 from 2020 (prior year). 5. September 25 Pay repairs and maintenance expenses, $11,400. 6. October 19 Purchase equipment for $6,400 cash. 7. December 30 Pay $1,100 cash dividends to stockholders.

7. Prepare an adjusted trial balance.
Answer #1

ADJUSTED TRIAL BALANCE

PARTICULARS DEBIT CREDIT

CASH (NOTE-1)    31400   

A/C RECEIVABLE (NOTE-2) 9400

SUPPLIES 2400

EQUIPMENT (NOTE-3) 24400

ACCUMILATED DEPRICIATION 6200

SALARIES PAYABLE (NOTE-4)    0

COMMON STOCK (NOTE-5) 23000

RETAINED EARNINGS(NOTE-6) 38400

TOTAL 67600 67600

NOTES TO ACCOUNTS

CASH OPENING BALANCE = 12000

CASH RECEIVED FOR SERVICE = 24600

COLLECTION OF A/C RECEIVABLE = 16400

ISSUE OF COMMON STOCK = 6000

PAYMENT OF SALARIES = (8700)

PAYMENT FOR REPAIRS = (11400)

PURCHASE OF EQUIPMENT = (6400)

PAYMENT OF DIVIDEND = (1100)

CLOSING BAL = 31400

NOTE-2 ADJUSTMENT OF A/C RECEIVABLE

A/C REceivabe opening balance = 6400

services provided on credit = 19400

a/c receivables collected = (16400)

CLOSING BAL 9400

NOTE-3 EQUIPMENT

EQUIPMENT OPENING BALANCE = 18000

PURCHASED DURING YEAR = 6400

CLOSING BAL 24400

NOTE-4 SALARIES PAYABLE

OPENING BALANCE = 8700

PAID DURING THE YEAR = (8700)

CLOSING BALANCE == 0

NOTE -5 COMMON STOCK

OPENING BALANCE = 17000

ISSUED DURING THE YEAR = 6000

CLOSING BAL 23000

NOTE- 6 RETAINED EARNINGS

OPENING BALANCE = 6900

PROFIT( REFER P&L) = 31500

CLOSING BAL = 38400

STAEMENT OF PROFIT AND LOSS ACCOUNT

TO REPAIRS = 11400 BY SERVICE =44000
TO DIVIDEND = 1100
TO NET PROFIT= 31500

NOTE: AS INFORMATION RELATED TO DEPRICIATION IS NOT SPECIFIED IN QUESTION HENCE DEPRICIATION IS NOT CHARGED

THANK YOU